OpenStudy (anonymous):

Can anyone help with calculus homework? I am a failing calculus learner that needs a lot of help with homework that is due today. Any help would be appreciated I need to get this class over with.

hero (hero):

Do you need "help" or do you need someone to do it for you?

OpenStudy (anonymous):

both lol

OpenStudy (anonymous):

what is that

hero (hero):

It's something that will help you understand a few things

OpenStudy (anonymous):

I participate when I ask questions to learn but I am not understanding calculus and I need help to better understand it

OpenStudy (anonymous):

we don't have any tutoring for this subject so I am looking for help is all

hero (hero):

http://openstudy.com/messages

hero (hero):

@larissa77 are you still here?

OpenStudy (anonymous):

yes

hero (hero):

Did you check your messages yet? What are you working on in Calculus?

OpenStudy (anonymous):

alot of things I am not sure what you are asking me. I went to the website and that isn't going to help me no one is online

hero (hero):

You don't know what topics you're working on in Calculus?

OpenStudy (anonymous):

derivative, tangent line, relative extrema, time, height function equation, rate of change. marginal profit, actual gain

OpenStudy (anonymous):

Here is what I am working on I will send you the homework assignment so you can see it. I really cant explain it when I am lost.

hero (hero):

The whole unit?

OpenStudy (anonymous):

what do you mean whole unit

hero (hero):

Have you tried watching some Khan Academy videos on these topics?

OpenStudy (anonymous):

I have tried everything to help me but its hard when I never got algebra and now calculus just confuses me more

OpenStudy (anonymous):

and the professor is only able to help for two hours that's for all her students at once and she had to help me finish last week's assignment

OpenStudy (anonymous):

so when you have nobody where do you turn too

OpenStudy (usukidoll):

My calc i professor told me that if you don't know your algebra, you will not do good in Calculus at all.

OpenStudy (amistre64):

pick one thing to focus on instead of trying to do 'everything' at once. pick one question/topic, and we can work thru it.

OpenStudy (anonymous):

like find the derivative of the following functions y=2x^2-5x+3x^-2+4x^-3 This is what I came up with dy/dx=d/dx (2x^2-5x+3x^-2+4x^-3 =d/dx (2x^2)-d/dx (5x) + d/dx (3x^-2) + d/dx (4x^-3) = 4x - 5x + -6x + -12x

OpenStudy (amistre64):

this one ... found it

OpenStudy (amistre64):

=d/dx (2x^2)-d/dx (5x) + d/dx (3x^-2) + d/dx (4x^-3) this part is fine, we can step it out further by pulling out the constants right? =2 d/dx (x^2) -5 d/dx (x) +3 d/dx (x^-2) +4 d/dx (x^-3) now its just a bunch of power rules

OpenStudy (amistre64):

the d/dx parts are best seen in the equation editor as: \[=2 \frac{d}{dx} (x^2) -5 \frac{d}{dx} (x) +3 \frac{d}{dx} (x^{-2}) +4 \frac{d}{dx} (x^{-3})\] \[=2 \frac{dx}{dx} (2x) -5 \frac{dx}{dx} (1) +3 \frac{dx}{dx} (-2x^{-3}) +4 \frac{dx}{dx} (-3x^{-4})\] and since dx/dx=1 \[=2 (2x) -5(1) +3 (-2x^{-3}) +4 (-3x^{-4})\]

OpenStudy (anonymous):

here's the next question f(t)= (5t^3 + t^2) (2t -4) here's what I got f (x) = (5t^3 + t^2) f' (x) = (15t^2 + 2t) t (x) = (2t -4) t' (x) = 2 f' (t) = (15t^2 + 2t) (2t -4) + 2(5t^3 + t^2)

OpenStudy (amistre64):

looks like youre trying for the product rule. youve named the parts badly, but the process seems fine

OpenStudy (amistre64):

f(t) = g(t) h(t) might be a better representation of it f' = g'h + gh' and you have g and h derived fine. do they want you to simplify any more? my teacher wasnt that concerned with the algebra afterwards as long as we could show that we knew how to work the derivative

OpenStudy (anonymous):

I will give you the file to the assignment so you can see what I am working on

OpenStudy (amistre64):

its all pretty basic. you seem to be doing alright so far. the next one would be simpler if you split the fraction; but they might want you to not do that and just work a quotient rule

OpenStudy (amistre64):

my times up, gotta go back to the house so im not going to be online ... post your workings and there are plenty of smart people around to chk your work and giv eyou some guidance :)

OpenStudy (anonymous):

ok